# Bulk photovoltaic effect in ferroelectric and antiferroelectric phases of antimony sulphoiodide investigated by means of ab-initio simulations

### v1

@@ Abstract
First-principles calculations are employed to study the ferroelectric properties and the bulk photovoltaic effect (BPVE) of antimony sulfur iodide (SbSI). The BPVE enables direct sunlight-to-electricity conversion in homogeneous materials and, in ferroelectric compounds, can be tuned via an electric field controlling the polarization. However, most ferroelectrics are oxides with large band gaps exceeding visible light energy, limiting their photovoltaic performance. Featuring a visible-range band gap, SbSI instead combines remarkable photovoltaic capabilities with a spin-textured band structure, coupling charge and spin degrees of freedom. The calculations predict ferroelectric and antiferroelectric phases with comparable band gaps but distinct Rashba-Dresselhaus and Weyl-like spin textures relevant for spintronics. The BPVE is driven by the linear and circular photogalvanic effects, exhibiting large photoconductivities under visible light. Furthermore, it can serve to identify the phase of the material, with the circular photogalvanic effect which also reflects spin-texture changes. With its multifunctional properties, SbSI emerges as a promising material candidate for efficient solar energy conversion and advanced electronics, with potential applications extending to spintronics.
@@
E-mail: silvia.picozzi@unimib.it
@@ 1 Introduction
The bulk photovoltaic effect (BPVE) converts solar power into electrical power within homogeneous systems [1, 2, 3, 4]. Unlike the conventional photovoltaic effect, which requires a built-in potential at material interfaces (e.g., p–n junctions) to separate the photogenerated charge carriers, the BPVE allows for simpler device structures where the entire bulk material contributes to the photocurrent generation. Furthermore, the BPVE can result in above-band-gap open circuits voltages, as carriers are collected before they thermalize, thereby surpassing the Shockley–Queisser limit which constrains conventional solar cells [5, 6].
The primary contributions to the BPVE are the shift and injection currents [7, 4, 8, 9], both second-order responses to the incident light field. The shift current [10] is an interband phenomenon, which can be interpreted as a real-space displacement of electrons upon light excitation [4, 11, 12]. In contrast, the injection current arises from asymmetries in the velocity distribution of excited carriers [1] due to non-reciprocal band structures [13] or processes like asymmetric carriers photoexcitation, recombination, and momentum relaxation [14]. Recently, either the shift current, the injection current, or both have been investigated experimentally [15, 16, 17, 18, 19, 20, 21, 22, 23, 24] as well as theoretically [25, 26, 27, 28] across numerous systems.
@@
Among the materials exhibiting both photogalvanic effects, ferroelectric (FE) compounds with spontaneous electrical polarization have attracted significant interest [1, 29, 30, 31, 32, 33], as the photocurrent direction can be controlled by reversing the polarization via an electric field. Despite that, the most studied FE materials, such as the perovskite oxides LiNbO3, BaTiO3 [29, 30], BiFeO3 [31] and PbTiO3 [30] have large band gaps greater than 3 eV, limiting their ability to capture sunlight to the ultraviolet (UV) range, which accounts for only about 3.5% of the total solar spectrum. For efficient photoconversion, ideal materials should possess narrower band gaps, strong absorption in the visible spectrum [34], along with long carrier lifetimes and diffusion length [35].
@@
Another promising alternative to oxide materials is antimony sulfoiodide (SbSI), which has a visible-range band gap [40], equal to approximately 1.9 eV. SbSI belongs to the class of multifunctional materials known as FE Rashba-Dresselhaus semiconductors [41, 42]. In these materials, strong spin-orbit coupling (SOC) is combined with ferroelectricity resulting in spin-textured electronic bands characterized by a large Rashba-Dresselhaus spin splitting that can be controlled by applying an electric field [42]. This effect can find applications in novel device concepts for spintronic applications [43]. Furthermore, the Rashba-Dresselhaus spin-texture is also predicted to lead to long photocarriers lifetimes [44, 45]. Consequently, FE Rashba-Dresselhaus semiconductors possess all the essential characteristics for efficient and tuneable photoconversion, which, moreover, can potentially be further integrated with electronic and spin transport functionalities.
@@
The discovery of ferroelectricity in SbSI dates back to 1962, with a reported spontaneous polarization of about 25 $\mu$C/cm2 at 0o C [46]. It was later found that SbSI also exhibits an antiferroelectric (AFE) phase in the temperature range 295 K $<\mathrm{T}<$ 410 K, above which the material transitions to the paraelectric phase. In the FE phase, SbSI exhibits high photoconductivity [47] and nonlinear optical behavior [48]. Moreover, the shift current has been experimentally measured [49, 50, 51], yielding an extrapolated photoconductivity of $\sim 10$ $\mu$A/V2, which is one of the largest reported values for a FE semiconductor. In contrast, the AFE phase, and its optical and photoconductive properties, have been far less studied, despite the AFE phase occurring near room temperature.
In this paper, we use first-principles calculations based on Density Functional Theory (DFT) to fill the knowledge gap regarding the different phases of SbSI, providing an accurate description of the BPVE. Specifically, we begin by comparing the FE and AFE structural and electronic properties, predicting that the phase transition also leads to a drastic modification of the band structure spin-texture, which could be of potential interest for spintronics. We then proceed with a detailed analysis of the BPVE, accurately estimating the shift and circular injection photoconductivities, and demonstrating that the latter is also related to the band structure spin-texture. Two main findings emerge from our study. First, the photocurrent can reach very large peak values within the visible energy range in both the FE and AFE phases, confirming the potential of the material for photovoltaic applications. Second, the direction of both shift and circular photocurrents changes as the system transitions from the FE to the AFE phase, due to the change in the crystal space group. Given that the FE-AFE phase transition occurs near room temperature and the photoresponses are experimentally accessible, we then propose that the BPVE can serve as a reliable probe of the phase transition. Moreover, we also show that measuring circular injection currents can track the material’s evolution from a Rashba-Dresselhaus spin character in the FE phase to a Dresselhaus-Weyl character in the AFE phase, revealing the intricate coupling between structural and spintronic properties.
@@
The paper is organized as follows. The computational details are reported in section 2. The results are presented in section 3: we describe the structural properties in subsection 3.1, estimate the macroscopic polarization in subsection 3.2, analyze the band structure in subsection 3.3, and the photoconductivities in subsection 3.4. Finally, in section 4, we summarize our conclusions.
@@ 2 Computational details
The linear and circular photogalvanic effects are described in terms of their photoconductivity tensors, using the approach developed by Azpiroz et al. [11] and Puente-Uriona et al. [12]. This approach is based on interpolation of the DFT band structure using maximally-localized Wannier functions [60, 61], as implemented in the Wannier90 package [62]. The $s$ and $p$-orbitals of Sb, S, and I ions are considered, resulting in a total of 96 bands when spin-orbit coupling (SOC) is included. For bulk systems, a 100 $\times$ 100 $\times$ 100 $k$-grid is used together with the adaptive broadening scheme described in Ref. [63] to cope with the Dirac delta functions. For the single-chain system, a 1 $\times$ 1 $\times$ 100 $k$-grid is employed, with a fixed broadening of 0.02 eV to better handle the van Hove singularities, following the method of Azpiroz et al. [11]. With these $k$-grids, the photoconductivities are well converged, and no significant differences in the results are observed when the grid density is further increased. SOC is included in all calculations of the band structure and photocurrents, while it is neglected in the geometry optimizations, on which SOC is expected to have no effect.
@@ 3.1 Structural properties
In the PE phase, both cations and anions are placed at the same $z$ coordinate. In contrast, in the FE phase, a cooperative displacement of Sb atoms results in a shift of the cation sub-lattice along the $c$-axis relative to the anion I and S sub-lattices, as shown in Fig. 1 (c). The Sb ions play a primary role by displacing more than the other atoms with respect to their centrosymmetric positions, breaking inversion symmetry and creating a polarization along the $c$-axis [65]. This behavior can be attributed to the presence of a Sb3+ lone pair, extending in the empty region between the chains and tilting along the $c$ axis in the FE phase so to minimize the Coulomb repulsion with the other bonding electrons [65]. The crystal structure in the FE phase remains orthorhombic as in the PE phase, but the space group changes to Pna21.
The crystal structure of AFE phase has not been studied in detail to date. Here, we propose a configuration, where the cations in one chain shift in one direction along the $c$-axis as in FE phase and the cations in adjacent chains move in the opposite direction [Fig. 1(d)]. As a result, each chain exhibits a polarization, but the polarizations of neighboring chains are oriented in opposite (antiparallel) directions, leading to a net zero polarization for the overall system. The crystal structure of this AFE phase is P212121.
@@
The FE distortion can not be reproduced by means of DFT within the standard GGA functional, because the structural optimization tends to converge to the PE structure, as already reported in previous studies [65]. To overcome this problem, we perform ionic relaxations using the hybrid functional HSE06 [57], by also taking into account the Van der Waals interaction among the chains. The HSE06 functional is known to effectively capture the delicate balance between electronic properties and lattice distortions [66, 67].
@@
In our HSE calculations, the FE configuration is found to be lower in energy than the PE and AFE configurations by 4.20 meV and 2.59 meV per formula unit, respectively; in other words, the AFE configuration is the second lowest in energy. Therefore, our results reproduce the stability ordering of the three phases observed in experiments. The optimized lattice parameters and atomic positions in the PE, FE and AFE phases are reported in detail in the Supporting Information (SI). The obtained values are in good agreement with experiments. For example, in the FE phase, our optimized lattice parameters are $a$=8.56 Å, $b$=10.10 Å and $c$=4.16 Å, while the experimental ones are respectively 8.49 Å, 10.10 Å and 4.16 Å [68]. The lattice parameters only slightly differ in the AFE and FE phases. Notably, in both these phases, the chains tend to elongate along the $c$-axis with respect to the PE geometry. In fact, we find that the $c$ lattice parameter increases from 4.10 Å in the PE phase to 4.16 Å, and 4.17 Å in the FE and AFE phases, respectively. This points to a “strain-polarization” coupling: while polarization develops, the $c$-axis elongates, in close similarity to what happens for standard ferroelectrics, such as P4mm BaTiO3. In the PE phase, the atom labeled Sb3 in Fig. 1(c) is positioned at the same $z$ coordinate as the I1 and S3 atoms. In contrast, in both the FE and AFE phases, Sb3 is displaced along the $c$ axis relative to I1 and S3 by $\Delta z_{\mathrm{Sb}_{3}-\mathrm{I}_{1}}=0.29$  and $\Delta z_{\mathrm{Sb}_{3}-\mathrm{S}_{3}}=0.21$. These values, which quantify the polar distortion, are slightly overestimated but in reasonable agreement with the experimental values for the FE phase, which are 0.21 Å and 0.17 Å[69]. This demonstrates the good performance of HSE calculations in predicting the FE structure of the material.
@@ 3.2 Electrical polarization
The macroscopic electrical polarization, $\mathbf{P}$, is calculated within the modern theory of polarization through the Berry phase method [59]. We build an adiabatic path from the optimized PE to FE and AFE structures by linearly interpolating both lattice parameters and atomic positions, labeled by a generalized coordinate $\lambda$, as depicted in Fig. 2. The parameter runs from -1 (AFE structure with space group P212121) through 0 (PE structure with Pnam space group), and up to 1 (FE structure with Pna21 space group). The resulting energy profile shows a characteristic (asymmetric) double-well shape, with an energy barrier of 4.20 meV per f.u. between the FE and PE phase, and a lower energy barrier of 2.59 meV between the AFE and the PE phase, as anticipated in the previous subsection.
@@
The macroscopic polarization in the AFE phase is zero. However, we can estimate the polarization of a single chain approximately by computing the tensor of the Born effective (or dynamical) charges (BEC) [70, 71], $Z^{*}_{ab}$ with $a,b=x,y,z$. The contribution of each atom to polarization can be defined as the product of the BEC and the displacements of the atom in all three Cartesian directions relative to their positions in the PE configuration [72]. The polarization of a single chain is then obtained by summing over the contributions of all the atoms in the chain. In Table 1, we provide the components of the BEC tensor calculated for the PE phase, computed using the GGA functional. We also note that there are no significant differences between the BEC values obtained with the GGA and HSE functionals.
For consistency, we use the BEC values from the PE phase to calculate the polarization in both the FE and AFE configurations. The BEC tensor for the FE phase is reported in the SI. As expected, the BEC components are positive for the cations and negative for the anions. The BEC values are large, deviating from the nominal ionic charges (namely, $+3$ for Sb, $-1$ for I, and $-2$ for S) indicating substantial charge redistribution across the bonds compared to an ideal ionic model, in particular along the $z$ direction. The highest BEC values are obtained for the Sb atoms, which play the crucial role of most ”active sites” in the material’s polarization, as discussed earlier.

### v2

@@ Abstract
We employ first-principles calculations to investigate the ferroelectric properties and the bulk photovoltaic effect (BPVE) of antimony sulfur iodide (SbSI). The BPVE enables direct sunlight-to-electricity conversion in homogeneous materials and, in ferroelectric compounds, can be tuned via an electric field controlling the polarization. However, most ferroelectrics are oxides with large band gaps exceeding the energy of visible light, thereby limiting their photovoltaic performance. SbSI, featuring a visible-range band gap, combines remarkable photovoltaic capabilities with a spin-textured band structure, coupling charge and spin degrees of freedom. Our calculations predict ferroelectric and antiferroelectric phases with comparable band gaps but distinct spin textures, relevant for spintronics applications. The BPVE is driven by the linear and circular photogalvanic effects, exhibiting high photoconductivities under visible light. Furthermore, it serves as a diagnostic tool to identify the material phase, with the circular photogalvanic effect reflecting spin texture changes. Thanks to its multifunctional properties, SbSI emerges as a promising candidate for solar energy conversion and advanced electronics, with potential applications extending to spintronics.
@@
E-mail: giuseppe.cuono@spin.cnr.it, subha.7491@gmail.com, andrea.droghetti@unive.it, silvia.picozzi@unimib.it
@@ 1 Introduction
The bulk photovoltaic effect (BPVE) enables the conversion of solar energy into electrical power within homogeneous systems [1, 2, 3, 4]. Unlike the conventional photovoltaic effect, which relies on a built-in potential at material interfaces (e.g., p–n junctions) to separate photogenerated charge carriers, the BPVE allows for simpler device architectures where the entire bulk material contributes to photocurrent generation. Furthermore, the BPVE has been reported to produce above-band-gap open-circuit voltages, as charge carriers can be collected before thermalization, potentially exceeding the Shockley–Queisser limit that constrains conventional solar cells [5, 6]. However this claim remains under debate, as other factors may limit energy conversion efficiency [7, 8].
The primary contributions to the BPVE are the shift and injection currents [9, 4, 10, 11], both second-order responses to the incident light field. The shift current [12] is an interband phenomenon, which can be interpreted as a real-space displacement of electrons upon light excitation [4, 13, 14]. In contrast, the injection current arises from asymmetries in the velocity distribution of excited carriers [1] due to non-reciprocal band structures [15, stavric2025] or processes like asymmetric carriers photoexcitation, recombination, and momentum relaxation [16]. Recently, either the shift current, the injection current, or both have been investigated experimentally [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] as well as theoretically [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41] across numerous systems.
@@
Among the materials exhibiting both photogalvanic effects, ferroelectric (FE) compounds with spontaneous electrical polarization have attracted significant interest [1, 42, 43, 44, 45, 46, 31], as the photocurrent direction can be controlled by reversing the polarization via an electric field. Despite that, the most studied FE materials, such as the perovskite oxides LiNbO3, BaTiO3 [42, 43], BiFeO3 [44] and PbTiO3 [43] have band gaps greater than 3 eV, limiting their ability to capture sunlight to the ultraviolet (UV) range, which accounts for only about 3.5% of the total solar spectrum. Ideal materials for studying the BPVE should possess narrower band gaps and strong absorption in the visible spectrum [47].
@@
Among the FE materials, antimony sulfoiodide (SbSI) is another promising alternative to oxides owing to its visible-range band gap [52], equal to approximately 1.9 eV. SbSI belongs to the class of multifunctional materials known as FE Rashba-Dresselhaus semiconductors [53, 54]. In these materials, strong spin-orbit coupling (SOC) is combined with ferroelectricity resulting in spin textured electronic bands characterized by a large Rashba-Dresselhaus spin splitting that can be controlled by applying an electric field [54]. This effect can find applications in novel device concepts for spintronic applications [55]. Thus, FE Rashba-Dresselhaus semiconductors possess the essential characteristics for the BPVE, which, moreover, can potentially be further integrated with electronic and spin transport functionalities.
@@
The discovery of ferroelectricity in SbSI dates back to 1962, with a reported spontaneous polarization of about 25 $\mu$C/cm2 at 0o C [56]. It was later found that SbSI also exhibits an antiferroelectric (AFE) phase in the temperature range 295 K $<\mathrm{T}<$ 410 K, above which the material transitions to the paraelectric phase [57, 58]. In the FE phase, SbSI exhibits high photoconductivity [59] and nonlinear optical behavior [60]. The shift current has been experimentally measured in several studies [61, 62, 63, 64], yielding an extrapolated photoconductivity of $\sim 10$ $\mu$A/V2 - one of the largest reported values for a FE semiconductor. While significant progress has been made in studying this phenomenon, a comprehensive theoretical analysis that derives all components of the photoconductivity tensor, evaluates their relative contributions across the visible spectrum, and explores the upper bounds of photocurrent generation remains to be fully developed. In addition, the AFE phase has been comparatively less explored, despite its stability near room temperature.
In this paper, we use first-principles calculations based on Density Functional Theory (DFT) to fill this knowledge gap. Specifically, we begin by comparing the FE and AFE structures, elucidating the mechanisms that stabilize these phases. We then demonstrate that the phase transition is accompanied by a significant modification of the spin texture in the band structure - a feature that may be of interest for spintronic applications. Finally, we analyze the BPVE, accurately computing both the shift and circular injection photoconductivities by deriving all symmetry-allowed tensor components and estimating their magnitudes across the visible energy range.
Three main findings emerge from our study. First, the photocurrent can reach very large peak values, significantly exceeding those observed in experiments to date. This points to untapped potential for enhancing the material’s photovoltaic performance. Second, we identify how the shift and circular photocurrents evolve across the FE-to-AFE transition. These insights can be potentially important for setting up measurements capable of distinguishing between the two phases. Third, we show how circular injection current measurements can track the change in spin texture from the FE phase to the AFE phase, revealing a rich interplay among structural, optical, and spintronic properties.
@@
The paper is organized as follows. The computational details are provided in Section 2. The results are presented in Section 3, where we describe the structural properties in Subsection 3.1, estimate the macroscopic polarization in Subsection 3.2, analyze the band structure in Subsection 3.3, and discuss the photoconductivities in Subsection 3.4. Finally, we summarize our conclusions in Section 4.
@@ 2 Computational details
The linear and circular photogalvanic effects are described in terms of their photoconductivity tensors, using the approach developed by Azpiroz et al. [13] and Puente-Uriona et al. [14]. This approach is based on interpolation of the DFT band structure using maximally-localized Wannier functions [73, 74], as implemented in the Wannier90 package [75]. The $s$ and $p$-orbitals of Sb, S, and I ions are considered, resulting in a total of 96 bands when spin-orbit coupling (SOC) is included. For bulk systems, a 100 $\times$ 100 $\times$ 100 $k$-grid is used together with the adaptive broadening scheme described in Ref. [76] to cope with the Dirac delta functions. For the single-chain system, a 1 $\times$ 1 $\times$ 100 $k$-grid is employed, with a fixed broadening of 0.02 eV to better handle the van Hove singularities, following the method of Azpiroz et al. [13]. With these $k$-grids, the photoconductivities are well converged, and no significant differences in the results are observed when the grid density is further increased. SOC is included in all calculations of the band structure and photocurrents, while it is neglected in the geometry optimizations, as it is not expected to significantly affect the structural parameters.
Symmetry mode analysis was performed using the ISODISTORT software [77, 78] to identify the atomic distortions associated with the AFE and FE phases, using the paraelectric (PE) Pnam phase as the reference structure.
@@ 3.1 Structural properties
We begin by presenting the structural and electronic properties of SbSI. While the FE phase has been previously investigated [79], the AFE phase has received comparatively little attention, and a comprehensive theoretical comparison between the two phases is still lacking. The results will also lay the groundwork for understanding the photoresponse.
@@
In the PE phase, both cations and anions are placed at the same $z$ coordinate. In contrast, in the FE phase, a cooperative displacement of Sb atoms results in a shift of the cation sub-lattice along the $c$-axis relative to the anion I and S sub-lattices, as shown in Fig. 1 (c). By performing a symmetry mode analysis, we identify this atomic distortion as the irreducible $\Gamma_{4}^{-}$ mode relative to the PE Pnam structure. The $\Gamma_{4}^{-}$ mode lowers the symmetry of the system to Pna21, while preserving the orthorhombic character. The Sb ions play a primary role by displacing more than the other atoms with respect to their centrosymmetric positions, breaking inversion symmetry and creating a polarization along the $c$-axis [79]. This behavior can be attributed to the presence of a Sb3+ lone pair, extending in the empty region between the chains and tilting along the $c$ axis in the FE phase so to minimize the Coulomb repulsion with the other bonding electrons [79].
The AFE phase crystallizes in the orthorhombic P212121 structure, as reported in experiments [57, 58]. Our symmetry mode analysis identifies the $\Gamma_{1}^{-}$ distortion relative to the Pnam PE phase as the driving mechanism [see Section S1 in the Supplemental Material (SM) [81]]. This $\Gamma_{1}^{-}$ mode corresponds to an antipolar distortion of the cations (along the $c$-axis) between adjacent atomic chains, as shown in Fig. 1(d). As a result, each chain exhibits polarization, but the polarizations of neighboring chains are oriented in opposite (antiparallel) directions, leading to a net zero polarization for the overall system. We note that, in addition to the $\Gamma_{1}^{-}$ mode, the fully relaxed P212121 structure also includes a $\Gamma_{1}^{+}$ atomic distortion and related $\eta_{\Gamma_{1}^{+}}$ strain, which, however, do not break the Pnam symmetry.
@@
The FE distortion can not be reproduced by means of DFT within the standard GGA functional because the structural optimization tends to converge to the PE structure, as already reported in previous studies [79]. To overcome this problem, we perform ionic relaxations using the hybrid functional HSE06 [70], by also taking into account the Van der Waals interaction among the chains. The HSE06 functional is known to effectively capture the delicate balance between electronic properties and lattice distortions [82, 83]. Similarly to the FE phase, the HSE functional is also found to be crucial for relaxing the AFE phase into the P212121 structure.
In our HSE calculations, the FE configuration is found to be lower in energy than the PE and AFE configurations by 4.20 meV and 2.59 meV per formula unit, respectively; in other words, the AFE configuration is the second lowest in energy. Therefore, our results reproduce the stability ordering of the three phases observed in experiments. Moreover, since the FE and AFE phases are close in energy, this suggests that the material is likely to exhibit antiferroelectricity at finite temperatures [84].
@@
The optimized lattice parameters and atomic positions in the PE, FE and AFE phases are reported in detail in the Section S2 of the SM. The obtained values are in good agreement with experiments. For example, in the FE phase, our optimized lattice parameters are $a=8.56$, $b=10.10$ and $c=4.16$, while the experimental ones are respectively 8.49 Å, 10.10 Å and 4.16 Å [85]. The lattice parameters only slightly differ in the AFE and FE phases. Notably, in both these phases, the chains tend to elongate along the $c$-axis with respect to the PE geometry. In fact, we find that the $c$ lattice parameter increases from 4.10 Å in the PE phase to 4.16 Å, and 4.17 Å in the FE and AFE phases, respectively. This points to a “strain-polarization” coupling: while polarization develops, the $c$-axis elongates, in close similarity to what happens for standard ferroelectrics, such as P4mm BaTiO3. In the PE phase, the atom labeled Sb3 in Fig. 1(c) is positioned at the same $z$ coordinate as the I1 and S3 atoms. In contrast, in both the FE and AFE phases, Sb3 is displaced along the $c$ axis relative to I1 and S3 by $\Delta z_{\mathrm{Sb}_{3}-\mathrm{I}_{1}}=0.29$  and $\Delta z_{\mathrm{Sb}_{3}-\mathrm{S}_{3}}=0.21$. These values, which quantify the polar distortion, are slightly overestimated but in reasonable agreement with the experimental values for the FE phase, which are 0.21 Å and 0.17 Å[86]. This demonstrates the good performance of HSE calculations in predicting the FE structure of the material.
@@ 3.2 Electrical polarization
[…]
@@
The macroscopic polarization in the AFE phase is zero. However, we can estimate the polarization of a single chain approximately by computing the tensor of the Born effective (or dynamical) charges (BEC) [87, 88], $Z^{*}_{ab}$ with $a,b=x,y,z$. The contribution of each atom to the polarization can be defined as the product of the BEC and the displacements of the atom in all three Cartesian directions relative to their positions in the PE configuration [89]. The polarization of a single chain is then obtained by summing over the contributions of all the atoms in the chain. In Table 1, we provide the components of the BEC tensor for the PE phase, computed using the GGA functional. We note that there are no significant differences between the BEC values obtained with the GGA and HSE functionals.
For consistency, we use the BEC values from the PE phase to calculate the polarization in both the FE and AFE configurations. The BEC tensor for the FE phase is reported in Section S4 the SM. As expected, the BEC components are positive for the cations and negative for the anions. The BEC values are large, deviating from the nominal ionic charges (namely, $+3$ for Sb, $-1$ for I, and $-2$ for S) indicating substantial charge redistribution across the bonds compared to an ideal ionic model, in particular along the $z$ direction. The highest BEC values are obtained for the Sb atoms, which play the crucial role of most “active sites” in the material’s polarization, as discussed earlier.